OpenStudy (anonymous):

Solve for x. x - 10 = -30

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

What number minus 10 equals -30?

OpenStudy (mathstudent55):

Since 10 is being subtracted from the left side, what do you need to do to the left side to undo that subtraction?

OpenStudy (anonymous):

subtract 10 from both sides?

OpenStudy (anonymous):

x=-20

OpenStudy (mathstudent55):

No. To undo an operation, you need to do the opposite operation. What is the opposite operation to subtraction?

OpenStudy (anonymous):

Oh you add 10 to both sides ok

OpenStudy (uri):

x - 10 = -30 x=-30+10 x=-20 No?

OpenStudy (mathstudent55):

That's it. Add 10 to both sides to undo the subtraction on the left side. This will give you just x on the left side and the answer on the right side.

OpenStudy (anonymous):

x=40

OpenStudy (mathstudent55):

No. The right side is _negative_ 30, not 30. What is -30 + 10?

OpenStudy (anonymous):

Its -20

OpenStudy (mathstudent55):

Great. Now let's check the solution. x - 10 = -30 We are saying that x = -20, so let's substitute -20 in for x: -20 - 10 = -30 -30 = -30 Since we have a true statement, x = -20 is correct.

OpenStudy (anonymous):

Ok thank you so much.

OpenStudy (mathstudent55):

wlcm

OpenStudy (uri):

So actually x=-20 is the answer..?

OpenStudy (mathstudent55):

Yes. Solution: x = -20

OpenStudy (uri):

Why did you go in so much depth? :P bored aye?

OpenStudy (uri):

@mathstudent55

OpenStudy (mathstudent55):

What?